OpenStudy (daniellelovee):

Sensors detect the concentration of mercaptan molecules in natural gas at various distances from a leaking source. At which location will a sensor detect the highest concentration of mercaptan?

OpenStudy (daniellelovee):

one-half meter in front of the source?

OpenStudy (daniellelovee):

@Somy

OpenStudy (somy):

hmmm it does sound right if you picture sensor as a wall if molecules are closer to the wall they are all detected while the ones at bigger distance from it - cannot be detected can they?

OpenStudy (daniellelovee):

nope

OpenStudy (daniellelovee):

alright thank you ^.^

OpenStudy (somy):

no problem :)

OpenStudy (daniellelovee):

can I ask one more and I'll stop bothering you lol

OpenStudy (daniellelovee):

You place a balloon in a closed chamber at standard temperature and pressure (STP). You double the chamber pressure. What happens to the balloon?

OpenStudy (somy):

lol its okay go ahead :)

OpenStudy (daniellelovee):

this are the options It expands to four times its volume. It stays the same volume. It doubles in volume. It decreases in volume by one-half.

OpenStudy (daniellelovee):

but I think it double in volume

OpenStudy (somy):

PV=nRT lets make nR as subject so that we get the P,V and T on one side nR= PV/T so initially its PV/T but then they doubled the P so 2PV/T

OpenStudy (somy):

now in the options they talk about volume so lets make volume as subject in case if you dont get it here V= nRT/P V=nRT/2P what do you see here

OpenStudy (somy):

initially nRT was divided by P and it was equal to V but now P is doubled! so nRT is divided by DOUBLE of P - in other words a bigger number so what will happen to V?

OpenStudy (daniellelovee):

is doubled :)

OpenStudy (somy):

no look its DIVIDED by INCREASED P value

OpenStudy (somy):

Ex. 2/2 = 1 2/4 = ?

OpenStudy (daniellelovee):

1/2

OpenStudy (somy):

bingo!

OpenStudy (daniellelovee):

so it decreased by a bigger value than itself therefore leading to an increase which means that A is the answer?

OpenStudy (daniellelovee):

sorry I meant D

OpenStudy (somy):

right! if you did mean D lol

OpenStudy (daniellelovee):

lol I did xD thank you! again xD

OpenStudy (somy):

you are welcome :))